The specifications for a certain kind of ribbon call for a mean breaking strength of 180 pounds. If five pieces of the ribbon (randomly selected from different rolls) have a mean breaking strength of 169.5 pounds with a standard deviation of 5.7 pounds, test the null hypothesis μ = 180 pounds against the alternative hypothesis μ < 180 pounds at the 0.01 level of significance. Assume that the population distribution is normal.

a) Find the p value

b) Test the hypothesis

c) Find the 99% C
Answer #1

\mu =180

\mu <180

\overline{X}=169.5,S=5.7,n=5

Significance level:\alpha =0.01

Test statistic: T=\frac{\overline{X}-\mu _0}{\frac{S}{\sqrt{n}}}=\frac{169.5-180}{\frac{5.7}{\sqrt{5}}}=-4.12

the degree of freedom = n-1=4

a) P value for the one-tailed test =0.0073

b) We can reject the null hypothesis as the P value is less than the significance level of the test.

Hence we have evidence that mean breaking strength is less than 180 pounds.

c)

For 99% confidence and df = 4

T_{\frac{\alpha }{2}}=4.6041

CI=\overline{X}\pm T_{\frac{\alpha }{2}}(\frac{S}{\sqrt{n}})=169.5\pm 4.6041(\frac{5.7}{\sqrt{5}})

CI=169.5\pm 11.74=(157.76,181.24)